OpenStudy (anonymous):

I'll give a medal if you help me :D Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 2, -4, and 1 + 3i

OpenStudy (anonymous):

from the zeros, make the factors then calculate the product of these factors and you will have your polynomial finally arrange the terms in decreasing power of variable.

OpenStudy (anonymous):

I tried but I can't seem to factor the 1+3i and the 1-3i

OpenStudy (anonymous):

(x-2)(x+4)[x-(1+3i)][x-(1-3i)] (x-2)(x+4)(x-1-3i)(x-1+3i) calculate this product

OpenStudy (anonymous):

how do you factor (x-1-3i)(x-1+3i)?

OpenStudy (anonymous):

you don't have to factor......you have to multiply them

OpenStudy (anonymous):

I meant multiply, sorry, do you multiply everything by the 2 1's or is 1-3i a single term? I keep getting confused on that part

OpenStudy (anonymous):

x, -1 and -3i are to be treated as independent of each other....

OpenStudy (anonymous):

(x-2)(x+4)[x-(1+3i)][x-(1-3i)] (x-2)(x+4)(x-1-3i)(x-1+3i) for first two brackets (x-2)(x+4) = (x^2 - 2x + 4x - 8) = (x^2 + 2x -8)

OpenStudy (anonymous):

is it x^4 + x^2 + 34x - 72?

OpenStudy (anonymous):

Have not calculated complete answer....hold on

OpenStudy (anonymous):

for next two brackets (x-1-3i)(x-1+3i) = (x^2 - x - 3ix - x + 1 + 3i + 3ix - 3i - 9 i^2) = (x^2 - 2x + 1 - 9*-1) = (x^2 - 2x + 10)

OpenStudy (anonymous):

So, (x-2)(x+4)[x-(1+3i)][x-(1-3i)] = (x^2 + 2x -8)(x^2 - 2x + 10) = (x^4 + 2x^3 - 8x^2 - 2x^3 - 4x^2 + 16x + 10x^2 + 20x - 80) = (x^4 - 2x^2 + 36x - 80)

OpenStudy (anonymous):

My answer is different from yours...............pls check if my calculations are correct .......

OpenStudy (anonymous):

I think you're right, I probably made a mistake in the multiplying :P

OpenStudy (anonymous):

ok.....so now you know how to handle imaginary roots.....

OpenStudy (anonymous):

yeah, thank you so much!

OpenStudy (anonymous):

u r welcome....☺